two angles are complementary. The measure of the larger angle is 5 times the measure of the smaller angle. Find the degree measure of the larger angle.
Answer by oberobic(2304)

Two angles: a & b.
They are complementary, so
a + b = 90
.
One angle is 5 times the other.
.
a = 5b
.
substitute a=5b
.
5b + b = 90
6b = 90
b = 15
.
a + b = 90
a = 90 -15
a = 75
.
Answer: The two angles are 15 and 75 degrees.
.
Done.